There are twice as many adult tickets sold as children's tickets. If adult tickets cost $8 and children's tickets cost $5 and if $4200 was collected, how many of each type were sold?
Answer by chriswen(106)

Let x be the amount of children's tickets sold.
Let 2x be the amount of adult ticket's sold.
...
8(2x)+5(x)=4200
16x+5x=4200
21x=4200
x=4200/21
x=200
2x=400
...
Therefore, 200 children's tickets were sold and 400 adult ticket's were sold.